I don't understand hoe to solve this word problem. It says King Kong and Godzilla are both busily destroying cities. The ratio of cities destroyed by King Kong tp those destroyed by Godzilla is 2:5. If the two together destroy a total of 35 cities, find how many cities each destroyed. Use the 5-step method. please include a check.
thank you

Found 2 solutions by jorel1380, chessace:
Answer by jorel1380(3719)

2n+5n=35
7n=35
n=5
King Kong destroys 2n cities, or 10 cities. Godzilla destroys 5n cities, or 25 cities. Together they destroy 35 cities. (How many times can you destroy Tokyo?)..

Answer by chessace(471)

Question is to find both..
K=cities destroyed by KK,
G=similar.
K:G = 2:5
Cross multiply:
5K=2G
G=(5/2)K
K+G=35= (1+5/2)K
(7/2)K=35
K=10
G=25
10+25=35
Not totally sure which are the 5 steps.